Using the equation:

C3H8 + 5O2 → 3CO2 + 4H2O


1. Calculate the moles of O2 needed to make 1.35 mol H2O.





2. Calculate the moles of C3H8 needed to make 13.4 mol CO2.




3. Calculate the moles of CO2 produced when 0.458 mol C3H8 reacts.




4. Calculate the the moles of H2O produced from 11.2 mol O2




5. Calculate the moles of CO2 produced when 3.86 mol O2 reacts. 




6. Calculate the moles of H2O produced when 9.54 mol C3H8 reacts.




7. Calculate the moles of C3H8 needed to produce 35.6 mol H2O.




8. Calculate the moles of O2 reacted when 3.57 mol of C3H8 combusts.




9. Calculate the moles of H2O produced when 2.68 mol CO2 is made.




10. Calculate the moles of CO2 produced when 0.876 mol C3H8 reacts. 

1
Expert's answer
2021-02-05T04:08:32-0500

In all these cases the proportions including reaction coefficients should be used.

1. n(O2)=5*1.35/4=1.6875(mol)

2. n(C3H8)=1*13.4/3=4.47(mol)

3. n(CO2)=3*0.458/1=1.374(mol)

4. n(H2O)=11.2*4/5=8.96(mol)

5. n(CO2)=3*3.86/5=2.316(mol)

6. n(H2O)=9.54*4/1=38.16(mol)

7. n(C3H8)=1*35.6/4=8.9(mol)

8. n(O2)=5*3.57/1=17.85(mol)

9. n(H2O)=4*2.68/3=3.57(mol)

10. n(CO2)=3*0.876/1=2.628(mol)